# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 1.2.2. Tug-of-war in PDE and probability
As [38] surveys, the game theory connection identified in [41] has attracted a lot of attention from PDE specialists. New boundary rules for $\epsilon$ tug-of-war led to more regular game value functions in [2]. Heavy-tailed moves connect to the infinity fractional Laplacian in [7]. A noisy version of the game has been considered, in which the counter makes a random displacement of magnitude $c\epsilon$ at the end of each turn. The $p$-Laplacian [37] interpolates the classical $p=2$ Laplacian operator and the infinity version, for $p\in(2,\infty)$. In [42], the value of the noisy game to shown to converge to a $p$-harmonic extension of boundary data, for $p$ suitably chosen as a function of $c$: the survey [35] takes this perspective as central. A variant of this game has been used to study $p$-Laplacian obstacle problems [36]. The abundant PDE connections of tug-of-war are reviewed in the book [8].
@@ 1.2.3. Weaving together the two research strands
As of 2025, [27] and [41] have both garnered over five hundred citations, with no article citing both until [22, 26]. Despite the thematic similarities and coincidence of names in the economists’ and mathematicians’ tug-of-war, the two veins of research appear to have developed quite independently for decades. The economists’ work treats much more developed random decision-rules for turn victory than the mathematicians’ trivial fair-coin (or constant-bias [40]) versions, but the mathematicians’ studies have a much richer geometric flavour. Weaving together the two strands is a very natural aim, but important differences should be noted: player-funded tug-of-war has a highly discrete aspect, with players even on long integer-interval gameboards committing significant resources only in a bounded window around a pivot or battlefield location; while in $\epsilon$-tug of war on Euclidean domains, individual turns have asymptotically no weight, so that analytic connections emerge (via PDE). In this regard, Brownian Boost brings the two perspectives together.
Recently, [26] introduced in the setting of gameboard $\mathbb{Z}$ an infinite-turn version of player-funded tug-of-war called the Trail of Lost Pennies. This article systematized aspects of the economists’ treatment by classifying and finding explicit formulas for all Markov-perfect Nash equilibria. It quantified the discouragement effect, proving that equilibria exist sometimes when incentives are unequal but also presenting clear numerical evidence that such equilibria are fragile: when players’ relative incentive differs from one by more than a quantity of order $10^{-4}$, equilibria cease to exist.
In the next subsection, we will introduce a two-parameter family of Trail of Lost Pennies games that generalize the example in [26]. Scaling suitably, in a fine-mesh high-noise limit, it is these games that will enable our rigorous study of $\rho$-Brownian Boost. Beyond playing this role, the new games a further allow us to test the robustness of the conclusions of [26] in a broader context. The research presented in this article was initiated and inspired by a comment offered by a referee of [26] who noted how the $p$-Laplacian arises by interpolating fair-coin tug-of-war with noise and asked, “if the two-player game was mixed with some probability $\alpha$ with a random walk, do the dynamics of the game change?” We write $\kappa=1-\alpha$, with Brownian Boost arising in the high-noise limit $\kappa\searrow 0$. And we introduce a second dimension of perturbation by modelling each turn on a Tullock contest of exponent $\rho\in(0,\infty)$. The two-dimensional family bears out important aspects of the battlefield and discouragement effects. We offer a comprehensive classification of Markov-perfect Nash equilibria in a broad swathe of the parameter space, finding surprising effects that warrant further study. We propose directions of inquiry for the discrete and stochastic-differential tug-of-war games in the hope that further study of such games might warrant the attention of analysts, economists and probabilists.
The introduction has five further subsections. In the next three, we specify the Trail of Lost Pennies $\mathrm{TLP}(\kappa,\rho)$; express its Nash equilibria in terms of a four-parameter $\mathbb{Z}$-indexed system of equations ABMN($\kappa,\rho$); and present the explicit solution of this system. In the penultimate subsection, we return to Brownian Boost, explaining how it is approximated by the scaled discrete games, and recording the principal conclusions, on stake-profiles and asymptotic gameplay, concerning the continuum game via this regularization, in Theorem 1.18. The final subsection reports our results concerning the discrete games $\mathrm{TLP}(\kappa,\rho)$, including the fixed-$(\kappa,\rho)$ asymptotics Theorem 1.21 and the implications of this result for the battlefield, stake asymmetry and discouragement effects.
@@ 1.3. The Trail of Lost Pennies
$C_{+}=\sum_{k=0}^{\infty}S_{+}(k)\,\,\,\,\textrm{and}\,\,\,\,C_{-}=\sum_{k=0}^{\infty}S_{-}(k)\,$, (3)
@@
The information permitted is a little limited, but in fact most of the strategies needed for our study make do with even less. For time-homogeneous Markov-perfect strategies, the only pertinent data in the record $X:\llbracket 0,k\rrbracket\to\mathbb{Z}$ available at the outset of the (k+1)(k+1)st turn is the present counter location $X(k)$. As in Subsection 1.1.3, we call any such strategy $S$, namely one whose value on every path is determined by the path’s terminal value, time-invariant; and write $\mathcal{S}_{0}$ for the space of these strategies. When Mina and Maxine play the respective elements of a time-invariant strategy pair $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$, we will abusively denote the pair $(b,a)$, for $a,b:\mathbb{Z}\to[0,\infty)$ given by
$\textrm{$a_{i}\,=\,S_{+}(\psi)$ \, and \, $b_{i}\,=\,S_{-}(\psi)$ \, for any $i\in\mathbb{N}$ and $\psi\in\Lambda_{i}$}\,.$ (5)
For $(S_{-},S_{+})\in\mathcal{S}^{2}$, the law of gameplay in $\mathrm{TLP}(\kappa,\rho)$ given $X(0)=\ell$ governed by the strategy pair $(S_{-},S_{+})$ will be denoted $\mathbb{P}^{\ell}_{S_{-},S_{+}}$, with $\mathbb{E}^{\ell}_{S_{-},S_{+}}[\cdot]$ the corresponding expectation. Note also that the usage $(S_{-},S_{+})\in\mathcal{S}^{2}$ entails a conflict where the stake offered under $S_{-}$ at the (k+1)(k+1)st turn, which is formally $S_{-}\big(X\big|_{\llbracket 0,k\rrbracket}\big)$, is referred to simply as $S_{-}(k)$ in (3). We will continue with the simpler usage in most instances since there is little prospect of confusion.
@@
1.4. Time-invariant Nash equilibria and ABMN($\kappa,\rho$) solutions
@@
For $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$, set $m_{i}=\mathbb{E}^{i}_{S_{-},S_{+}}[P_{+}]$ and $n_{i}=\mathbb{E}^{i}_{S_{-},S_{+}}[P_{-}]$ for $i\in\mathbb{Z}$. The values $a_{i}$ and $b_{i}$ are determined by (5). Thus to each time-invariant strategy pair $(S_{-},S_{+})$ we associate a quadruple $(a,b,m,n):\mathbb{Z}\to[0,\infty)^{2}\times\mathbb{R}^{2}$, and conversely any such quadruple determines $(S_{-},S_{+})$.
@@
$\lim_{k\to\infty}m_{-k}=m_{-\infty}\,\,\,,\,\,\,\lim_{k\to\infty}m_{k}=m_{\infty}\,\,\,,\,\,\,\lim_{k\to\infty}n_{-k}=n_{-\infty}\,\,\,\,\textrm{and}\,\,\,\,\lim_{k\to\infty}n_{k}=n_{\infty}\,.$ (6)
@@
$m_{-\infty}<m_{\infty}\,\,\,\,\textrm{and}\,\,\,\,n_{\infty}<n_{-\infty}\,.$ (7)
The next result states the basic relaionship between the trail game and ABMN: a time-invariant strategy pair is a Nash equilibrium if and only if it is the $(b,a)$-projection of an element of ABMN($\kappa,\rho$). Note that the assertion is made only under the condition that $\rho\leq 1$.
@@
Let $(\kappa,\rho)\in(0,1]^{2}$, and let $(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})\in\mathbb{R}^{4}$ satisfy (7).
@@
$W\,=\,\big\{(\kappa,\rho)\in(0,1]\times(0,\infty):\rho^{2}\kappa\leq 1\big\}\,.$ (8)
The hypothesis $(\kappa,\rho)\in W$ will be recalled from time to time in our study of ABMN($\kappa,\rho$), but in fact it is always in force.
@@
1.5. Explicit ABMN solutions
1.5.1. Ingredients for solving ABMN
@@
$\gamma(\kappa,\rho,\beta)$ $=$ $\frac{(1-\kappa)\beta^{2\rho}+2(1-\rho\kappa)\beta^{\rho}+1+\kappa}{2(1+\beta^{\rho})^{2}}\,$, (9)
@@
$\phi_{0}(\kappa,\rho,\beta)$ $=$ $\frac{\beta\Big((1-\kappa)\beta^{2\rho}+2(1+\kappa\rho)\beta^{\rho}+1+\kappa\Big)}{(1-\kappa)\beta^{2\rho}+2(1-\kappa\rho)\beta^{\rho}+1+\kappa}$ (10)
@@
$\phi_{1}(\kappa,\rho,\beta)\,=\,\frac{\beta\Big((1+\kappa)\beta^{2\rho}+2(1-\kappa\rho)\beta^{\rho}+1-\kappa\Big)}{(1+\kappa)\beta^{2\rho}+2(1+\rho\kappa)\beta^{\rho}+1-\kappa}\,.$ (11)
The four functions are positive, because our minimal hypothesis, that $(\kappa,\rho)$ belongs to the set $W$ specified in (8), implies that $\kappa$ and $\kappa\rho$ are at most one, and every displayed coefficient is then non-negative.
@@
    As we will show in Lemma 2.3, $\phi_{0}(\kappa,\rho,\cdot)$ and $\phi_{1}(\kappa,\rho,\cdot)$ are increasing bijections on $(0,\infty)$. Consequently, the map that sends $\phi_{0}\in(0,\infty)$ to $\phi_{1}$ is well defined. We label this function $s:(0,\infty)\to(0,\infty)$. Thus, for any given $x\in(0,\infty)$, $s(x)=\phi_{1}(\kappa,\rho,\beta)$ for the unique value of $\beta\in(0,\infty)$ for which $\phi_{0}(\kappa,\rho,\beta)=x$.
@@ 1.5.2. The solution formulas
$\textrm{$(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})\in\mathbb{R}^{4}$ satisfies $m_{-\infty}<m_{\infty}$ and $n_{\infty}<n_{-\infty}$}\,.$ (12)
@@
Given the four parameters in (12) and the three noted symmetries, we may expect the reduced solution space to be parametrized by one free parameter. What is a natural choice for this? We propose two, one local, the other global. For $(a,b,m,n)\in$ ABMN($\kappa,\rho$), the local choice is the central ratio $\rm CenRatio$}, which we set to be $\frac{n_{0,-1}}{m_{-1,0}}$. The global choice is the solution’s Mina margin which, recall, is defined to be $\tfrac{n_{\infty,-\infty}}{m_{-\infty,\infty}}$, or $\tfrac{n_{-\infty}}{m_{\infty}}$ given our assumptions. The local choice is useful for describing explicit formulas for solutions. The global choice is less useful as a parameter, because it does not bijectively index solutions up to symmetry, but this global statistic is important for understanding the game-theoretic consequences of the form of the solutions.
@@
$m^{\rm def}_{-\infty}(x)=0\,,\,\,\,\,\textrm{and}\,\,\,\,m^{\rm def}_{k+1}(x)-m^{\rm def}_{k}(x)\,=\,\kappa\prod_{i=0}^{k}\big(c_{i}(x)-1\big)\,\,\,\textrm{for $k\in\mathbb{Z}$}\,$, (13)
@@
We distinguish two choices of currency revaluation for solutions with central ratio $x$. The default solution has $\mu=1$. The other choice is $\mu=m^{\rm def}_{\infty}(x)^{-1}$ where $m^{\rm def}_{\infty}(x)=\kappa\sum_{k\in\mathbb{Z}}\prod_{i=0}^{k}\big(c_{i}(x)-1\big)$ is Maxine’s default prize. This solution $(a,b,m,n)$ is sometimes convenient (and we label it next), since satisfies the simple boundary condition $(m_{-\infty},n_{\infty})=(0,1)$ and $m_{\infty}=1$.
@@
Remark. The representation of solutions in Theorem 1.16 is governed by orbits of the $(\kappa,\rho)$-parameterised map $s:(0,\infty)\to(0,\infty)$. For generic $(\kappa,\rho)\in(0,1]^{2}$, there is no explicit form for $s:\phi_{0}\mapsto\phi_{1}$. In the case $(\kappa,\rho)=(1,1)$ analysed in [26], $\phi_{0}=\beta(2\beta+1)$ and $\phi_{1}=\beta^{2}/(\beta+2)$. Since $\phi_{0}$ and $\phi_{1}$ appear linearly in coefficients of quadratic equations in the $\beta$-variable, $s$ has an explicit form, given in [26, Definition $2.18$]. When $\kappa\in(0,1)$ and $\rho=1$, $\phi_{0}$ and $\phi_{1}$ appear in coefficients of cubic equations in $\beta$, and $s$ may be expressed as a rational function of the unique positive root of a cubic polynomial. The generic inexplicitness of $s$ may disconcert at first, but its implications for this study have been limited to the use of approximate root solving in the numerical investigation of ABMN($\kappa,\rho$) elements.

### v4

@@ 1.2.2. Tug-of-war in PDE and probability
As [38] surveys, the game theory connection identified in [43] has attracted a lot of attention from PDE specialists. New boundary rules for $\epsilon$ tug-of-war led to more regular game value functions in [2]. Heavy-tailed moves connect to the infinity fractional Laplacian in [6]. A noisy version of the game has been considered, in which the counter makes a random displacement of magnitude $c\epsilon$ at the end of each turn. The $p$-Laplacian [37] interpolates, as $p$ ranges over $(1,\infty)$, between the total-variation ($p=1$) operator—whose level-set evolution is motion by mean curvature; see [31] for a game-theoretic approach—and the infinity Laplacian obtained in the high-$p$ limit. In [44], the value of the noisy game is shown to converge to a $p$-harmonic extension of boundary data, for $p$ suitably chosen as a function of $c$: the survey [35] takes this perspective as central. A variant of this game has been used to study $p$-Laplacian obstacle problems [36]. The abundant PDE connections of tug-of-war are reviewed in the book [7].
@@ 1.2.3. Weaving together the two research strands
As of 2025, [25] and [43] have both garnered over five hundred citations, with no article citing both until [20, 24]. Despite the thematic similarities and coincidence of names in the economists’ and mathematicians’ tug-of-war, the two veins of research appear to have developed quite independently for decades. The economists’ work treats much more developed random decision rules for turn victory than the mathematicians’ trivial fair-coin (or constant-bias [42]) versions, but the mathematicians’ studies have a much richer geometric flavour.
Weaving together the two strands is therefore a natural aim, but one should note important structural distinctions, between or within the strands. One concerns the asymptotic role of individual turns. Player-funded tug-of-war has an intrinsically discrete character, with players, even on long integer-interval gameboards, committing significant resources only in a bounded window around a pivot or battlefield location. By contrast, in $\varepsilon$-tug-of-war on Euclidean domains, individual turns have asymptotically negligible weight, and analytic connections emerge, via PDE.
A second distinction, for the stake-governed games, concerns the source of funding. In player-funded games, strategic expenditure is endogenous, while in budget-allocated games the spending constraint is imposed by the game design. Brownian Boost occupies an intermediate position. It is player-funded, yet shares the asymptotic feature that individual actions carry vanishing weight; its analytic structure is governed by coupled ODE rather than PDE.
When budgets are allocated and play takes place on a fine mesh in a Euclidean domain, the infinity-harmonic structure of value functions identified in discrete settings in [20] may plausibly give rise to PDE connections, with the optimal stake proportion (shared by the players) determined formulaically by the associated infinity-harmonic value function.
In this sense, the present work and [20] outline two coherent routes by which economically motivated contest models may merit the attention of analysts and probabilists.
@@ 1.3. The Trail of Lost Pennies
In three subsections, we specify the discrete games; derive a coupled system of four equations indexed by $\mathbb{Z}$ that govern equilibrium values and stakes; and present explicit solution formulas for this system.
1.3.1. Specifying the discrete games
@@
$C_{+}=\sum_{k=0}^{\infty}S_{+}(k)\,\,\,\,\textrm{and}\,\,\,\,C_{-}=\sum_{k=0}^{\infty}S_{-}(k)\,$, (1.2)
@@
The information permitted is a little limited, but in fact most of the strategies needed for our study make do with even less. For time-homogeneous Markov-perfect strategies, the only pertinent data in the record $X:\llbracket 0,k\rrbracket\to\mathbb{Z}$ available at the outset of the (k+1)(k+1)st turn is the present counter location $X(k)$. As in Section 1.4.2, we call any such strategy $S$, namely one whose value on every path is determined by the path’s terminal value, time-invariant; and write $\mathcal{S}_{0}$ for the space of these strategies. When Mina and Maxine play the respective elements of a time-invariant strategy pair $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$, we will abusively denote the pair $(b,a)$, for $a,b:\mathbb{Z}\to[0,\infty)$ given by
$\textrm{$a_{i}\,=\,S_{+}(\psi)$ \, and \, $b_{i}\,=\,S_{-}(\psi)$ \, for any $i\in\mathbb{N}$ and $\psi\in\Lambda_{i}$}\,.$ (1.4)
For $(S_{-},S_{+})\in\mathcal{S}^{2}$, the law of gameplay in $\mathrm{TLP}(\kappa,\rho)$ given $X(0)=\ell$ governed by the strategy pair $(S_{-},S_{+})$ will be denoted $\mathbb{P}^{\ell}_{S_{-},S_{+}}$, with $\mathbb{E}^{\ell}_{S_{-},S_{+}}[\cdot]$
the corresponding expectation. Note also that the usage $(S_{-},S_{+})\in\mathcal{S}^{2}$ entails a conflict where the stake offered under $S_{-}$ at the (k+1)(k+1)st turn, which is formally $S_{-}\big(X\big|_{\llbracket 0,k\rrbracket}\big)$, is referred to simply as $S_{-}(k)$ in (1.2). We will continue with the simpler usage in most instances since there is little prospect of confusion.
@@
1.3.2. Time-invariant Nash equilibria and ABMN($\kappa,\rho$) solutions
@@
For $(S_{-},S_{+})\in\mathcal{S}_{0}^{2}$, set $m_{i}=\mathbb{E}^{i}_{S_{-},S_{+}}[P_{+}]$ and $n_{i}=\mathbb{E}^{i}_{S_{-},S_{+}}[P_{-}]$ for $i\in\mathbb{Z}$. The values $a_{i}$ and $b_{i}$ are determined by (1.4). Thus to each time-invariant strategy pair $(S_{-},S_{+})$ we associate a quadruple $(a,b,m,n):\mathbb{Z}\to[0,\infty)^{2}\times\mathbb{R}^{2}$, and conversely any such quadruple determines $(S_{-},S_{+})$.
@@
$\lim_{k\to\infty}m_{-k}=m_{-\infty}\,\,\,,\,\,\,\lim_{k\to\infty}m_{k}=m_{\infty}\,\,\,,\,\,\,\lim_{k\to\infty}n_{-k}=n_{-\infty}\,\,\,\,\textrm{and}\,\,\,\,\lim_{k\to\infty}n_{k}=n_{\infty}\,.$ (1.5)
@@
$m_{-\infty}<m_{\infty}\,\,\,\,\textrm{and}\,\,\,\,n_{\infty}<n_{-\infty}\,.$ (1.6)
The next result states the basic relationship between the trail game and ABMN: a time-invariant strategy pair is a Nash equilibrium if and only if it is the $(b,a)$-projection of an element of ABMN($\kappa,\rho$). Note that the assertion is made only under the condition that $\rho\leq 1$.
@@
Let $(\kappa,\rho)\in(0,1]^{2}$, and let $(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})\in\mathbb{R}^{4}$ satisfy (1.6).
@@
$W\,=\,\big\{(\kappa,\rho)\in(0,1]\times(0,\infty):\rho^{2}\kappa\leq 1\big\}\,.$ (1.7)
The hypothesis $(\kappa,\rho)\in W$ will be recalled from time to time in our study of ABMN($\kappa,\rho$), but in practice it is almost always in force.
@@
1.3.3. Explicit ABMN solutions
Ingredients for solving ABMN
@@ 1.3.3. Explicit ABMN solutions
$\gamma(\kappa,\rho,\beta)$ $=$ $\frac{(1-\kappa)\beta^{2\rho}+2(1-\rho\kappa)\beta^{\rho}+1+\kappa}{2(1+\beta^{\rho})^{2}}\,$, (1.8)
@@
$\phi_{0}(\kappa,\rho,\beta)$ $=$ $\frac{\beta\Big((1-\kappa)\beta^{2\rho}+2(1+\kappa\rho)\beta^{\rho}+1+\kappa\Big)}{(1-\kappa)\beta^{2\rho}+2(1-\kappa\rho)\beta^{\rho}+1+\kappa}$ (1.9)
@@
$\phi_{1}(\kappa,\rho,\beta)\,=\,\frac{\beta\Big((1+\kappa)\beta^{2\rho}+2(1-\kappa\rho)\beta^{\rho}+1-\kappa\Big)}{(1+\kappa)\beta^{2\rho}+2(1+\rho\kappa)\beta^{\rho}+1-\kappa}\,.$ (1.10)
The four functions are positive, because our minimal hypothesis, that $(\kappa,\rho)$ belongs to the set $W$ specified in (1.7), implies that $\kappa$ and $\kappa\rho$ are at most one, and every displayed coefficient is then non-negative.
@@
    As we will show in Lemma 2.1, $\phi_{0}(\kappa,\rho,\bullet)$ and $\phi_{1}(\kappa,\rho,\bullet)$ are increasing bijections on $(0,\infty)$. Consequently, the map that sends $\phi_{0}\in(0,\infty)$ to $\phi_{1}$ is well defined. We label this function $s:(0,\infty)\to(0,\infty)$. Thus, for any given $x\in(0,\infty)$, $s(x)=\phi_{1}(\kappa,\rho,\beta)$ for the unique value of $\beta\in(0,\infty)$ for which $\phi_{0}(\kappa,\rho,\beta)=x$.
@@
$\textrm{$(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})\in\mathbb{R}^{4}$ satisfies $m_{-\infty}<m_{\infty}$ and $n_{\infty}<n_{-\infty}$}\,.$ (1.11)
@@
Given the four parameters in (1.11) and the three noted symmetries, we may expect the reduced solution space to be parametrized by one free parameter. What is a natural choice for this? We propose two, one local, the other global. For $(a,b,m,n)\in$ ABMN($\kappa,\rho$), the local choice is the central ratio $\rm CenRatio$}, which we set to be $\frac{n_{0,-1}}{m_{-1,0}}$. The global choice is the solution’s Mina margin which, recall, is defined to be $\tfrac{n_{\infty,-\infty}}{m_{-\infty,\infty}}$, or $\tfrac{n_{-\infty}}{m_{\infty}}$ given our assumptions. Both of these parameters offer measures of relative incentive: the central ratio expresses the average gain involved for Mina in winning a turn at which the counter starts at the origin relative to the average loss for Maxine due to the same outcome; while the Mina margin measures the terminal reward for Mina relative to this reward for Maxine. The local choice is useful for describing explicit formulas for solutions. The global choice is less useful as a parameter, because it does not bijectively index solutions up to symmetry, but this global statistic is important for understanding the game-theoretic consequences of the form of the solutions.
@@
$m^{\rm def}_{-\infty}(x)=0\,,\,\,\,\,\textrm{and}\,\,\,\,m^{\rm def}_{k+1}(x)-m^{\rm def}_{k}(x)\,=\,\kappa\prod_{i=0}^{k}\big(c_{i}(x)-1\big)\,\,\,\textrm{for $k\in\mathbb{Z}$}\,$, (1.12)
@@
We distinguish two choices of currency revaluation for solutions with central ratio $x$. The default solution has $\mu=1$. The other choice is $\mu=m^{\rm def}_{\infty}(x)^{-1}$ where $m^{\rm def}_{\infty}(x)=\kappa\sum_{k\in\mathbb{Z}}\prod_{i=0}^{k}\big(c_{i}(x)-1\big)$ is Maxine’s default prize. This solution $(a,b,m,n)$ satisfies the boundary condition $(m_{-\infty},n_{\infty})=(0,1)$ and $m_{\infty}=1$, which is a natural specification for game data. We label the solution accordingly.
@@
Remark. The representation of solutions in Theorem 1.11 is governed by orbits of the $(\kappa,\rho)$-parameterised map $s:(0,\infty)\to(0,\infty)$. For generic $(\kappa,\rho)\in(0,1]^{2}$, there is no explicit form for $s:\phi_{0}\mapsto\phi_{1}$. In the case $(\kappa,\rho)=(1,1)$ analysed in [24], $\phi_{0}=\beta(2\beta+1)$ and $\phi_{1}=\beta^{2}/(\beta+2)$. Since $\phi_{0}$ and $\phi_{1}$ appear linearly in coefficients of quadratic equations in the $\beta$-variable, $s$ has an explicit form, given in [24, Definition $2.18$]. When $\kappa\in(0,1)$ and $\rho=1$, $\phi_{0}$ and $\phi_{1}$ appear in coefficients of cubic equations in $\beta$, and $s$ may be expressed as a rational function of the unique positive root of a cubic polynomial. The generic inexplicitness of $s$ may disconcert at first, but its implications for this study have been limited to the use of approximate root solving in the numerical investigation of ABMN($\kappa,\rho$) elements.